OpenStudy (anonymous):

A deck of cards has three yellow cards and five green cards. One green card is drawn from the deck and not replaced. One more card is randomly drawn. Find the probability that the second card is yellow. three yellow,five green

OpenStudy (saifoo.khan):

Something.

OpenStudy (amistre64):

8! ---- is the sample space, but thats prolly useles info 3!5!

OpenStudy (anonymous):

the probability the second is yellow is P(first green, second yellow) + P(first yellow, second yellow), or 5/8 * 3/7 + 3/8 * 2/7 = 15/56 + 6/56 = 21/56

myininaya (myininaya):

i thought it would be 3/7

OpenStudy (llort):

is the first card replaced?

OpenStudy (anonymous):

but the first card chosen and not replaced might have been yellow.

myininaya (myininaya):

it said it was green

OpenStudy (llort):

oh it isn't, then it is 3/7

OpenStudy (anonymous):

my answer above is if you don't know the color of the first card drawn.

OpenStudy (anonymous):

oops.